### v1

Abstract
Atypical mitoses mark a deviation in the cell division process that can be an independent prognostically relevant marker for tumor malignancy. However, their identification remains challenging due to low prevalence, at times subtle morphological differences from normal mitoses, low inter-rater agreement among pathologists, and class imbalance in datasets. Building on the Atypical Mitosis dataset for Breast Cancer (AMi-Br), this study presents a comprehensive benchmark comparing deep learning approaches for automated atypical mitotic figure (AMF) classification, including baseline models, foundation models with linear probing, and foundation models fine-tuned with low-rank adaptation (LoRA). For rigorous evaluation, we further introduce two new hold-out AMF datasets - AtNorM-Br, a dataset of mitoses from the The TCGA breast cancer cohort, and AtNorM-MD, a multi-domain dataset of mitoses from the MIDOG++ training set. We found average balanced accuracy values of up to 0.8135, 0.7696, and 0.7705 on the in-domain AMi-Br and the out-of-domain AtNorm-Br and AtNorM-MD datasets, respectively, with the results being particularly good for LoRA-based adaptation of the Virchow-line of foundation models. Our work shows that atypical mitosis classification, while being a challenging problem, can be effectively addressed through the use of recent advances in transfer learning and model fine-tuning techniques. We make available all code and data used in this paper in this github repository: https://github.com/DeepMicroscopy/AMi-Br_Benchmark.
keywords:
Atypical Mitoses, Deep Learning, Foundation Models, Classification, Benchmarking, Histopathology, low-rank adaptation
volume: 3
1 Introduction
Figure 1: Image samples showing atypical and normal mitotic figures from all three datasets included in this study
Mitosis is the process where a cell replicates its genetic material (DNA) and then divides into two identical daughter cells. This is how cells grow, repair damaged tissues, and replace old or worn-out cells. The cell in the process of dividing to create two new cells may be observed microscopically as a mitotic figure (MF). Quantifying tumor proliferation is essential for assessing many human and animal cancers, including human breast cancer (Fitzgibbons and Connolly, 2023; Kiupel et al., 2011; Louis et al., 2016). A key metric in this assessment is the mitotic count, which reflects the number of cells in mitosis and provides valuable information about the tumor’s growth rate and, consequently, its histological grade. Elevated mitotic counts are associated with an increased risk of distant metastasis and reduced patient survival (Medri et al., 2003). Identification of MFs largely remains a manual task, making it highly subjective and time intensive. This task is furthermore subject to a high inter-rater variability, complicating the creation of datasets used for automatic detection (Meyer et al., 2005, 2009; Wilm et al., 2021). Deep learning models have been extensively used in tasks like mitosis detection and classification (Aubreville et al., 2024; Veta et al., 2019; Aubreville et al., 2023b), and they offer a promising outlook on the problem.
In tumor samples, we find two variations of MFs: normal and atypical. Normal MFs follow the typical, orderly process of cell division, including stages like prometaphase, metaphase, anaphase, and telophase, where chromosomes align and separate to produce two identical daughter cells Donovan et al. (2021). An atypical mitotic figure (AMF), on the other hand, is a cell undergoing an abnormal cell division process characterized by errors in chromosome segregation or other morphological irregularities, often seen in cancer cells (Donovan et al., 2021). The rate of AMFs has been correlated with poorer prognosis for breast cancer (Ohashi et al., 2018; Lashen et al., 2022) and other tumor types (Jin et al., 2007; Kalatova et al., 2015; Bertram et al., 2023; Matsuda et al., 2016). Recently, Jahanifar et al. (2025) have confirmed the prognostic relevance of the rate and density of AMFs across a wide range of tumors in a large-scale evaluation.
Yet, the identification of AMF is complicated by a low inter-rater agreement for classifying the mitoses into normal or atypical classes, due to challenging and irregular morphologies (Bertram et al., 2023; Aubreville et al., 2023a). Deep learning models have the potential to uniformly detect and classify mitoses, enable reproducibility, and reduce the workload of pathologists in this regard. Additionally, automatic, deep learning-based analysis allows for the large-scale, data-driven identification of prognostically relevant markers, such as the ratio of AMFs on the whole slide (Jahanifar et al., 2025).
There have been very few studies investigating deep learning for AMFs (Aubreville et al., 2023a; Fick et al., 2024; Bertram et al., 2025) The classification of atypical vs. normal MFs is complicated by the overlapping morphologies of the normal and atypical mitotic figures (Donovan et al., 2021), and the low frequency of AMF, resulting in class imbalance. Although this is a prognostically relevant problem, there is currently a lack of standardized benchmarks, which makes it difficult to compare approaches, hence it is crucial to develop suitable datasets and establish strong baselines that reflect the unique challenges of this task. Recently, foundation models, which are mostly vision transformer (ViT)-based models trained on millions of histopathology slides using self supervised learning, have gained popularity in medical imaging tasks like classification. These models are typically employed either via linear probing, where a lightweight classifier is trained on top of a frozen backbone, or through parameter-efficient fine-tuning techniques like Low Rank Adaptation (LoRA) (Hu et al., 2022). But there is a lack of comprehensive studies on using these methods for atypical vs. normal mitoses classification.
The contribution of this work is two-fold: First, we present three datasets, centered around the identification of AMFs, and secondly, we provide a comprehensive benchmark comparing widely used deep learning architectures, foundation models with linear probing, and foundation models fine-tuned with LoRA for the task of atypical versus normal mitosis classification.
2 Datasets
We present and use three distinct datasets in this paper - AMi-Br, AtNorM-Br and AtNorM-MD, as described below. This work is an extension of a previous conference contribution by our group (Bertram et al., 2025), which already introduced the first of those datasets (AMi-Br) in brief.
Being the largest scale dataset of the set, we use data from the AMI-Br dataset for training in all evaluations. This allows us to use the other two – newly introduced – datasets as independent hold out sets. It has been shown that data distribution shifts (domain shift) impede the performance of deep learning models significantly (Stacke et al., 2020; Aubreville et al., 2021). With these datasets, we can expect varying degrees of domain shifts, as they stem from different sources (AtNorm-Br) or even tumor types (AtNorm-MD), which allows us to investigate generalization. The statistics of the dataset, including total number of atypical and normal mitotic figures in each, annotation type, source dataset(s), species involved etc. are summarized in Table 1.
Table 1: Dataset statistics including sample counts, class balance, annotation details, and domain coverage.
Property AMi-Br AtNorM-MD AtNorM-Br
Total MFs 3,720 2,107 746
Atypical 832 219 128
Normal 2,888 1,888 618
Atypical Rate (%) 22.37% 10.39% 17.16%
Annotation Type 3-expert vote 5-expert vote Single expert
Expert Agreement 78.2% 69.6%
Source Dataset(s) TUPAC16, MIDOG21 MIDOG++ TCGA (BRCA)
Species Human Human + Canine Human
2.1 AMi-Br
The first dataset, AMi-Br (Bertram et al., 2025) consists of a total of 3,720 MFs from human breast cancer, of which 1,999 MFs are from the TUPAC16 (Veta et al., 2019) alternative label set (Bertram et al., 2020) and 1,721 MFs are from the MIDOG21 training dataset (Aubreville et al., 2023b). The MFs were classified into one of two classes — atypical or normal — by three expert pathologists. According to the result of the majority vote between the three experts, 832 MFs were found to be atypical, while the remaining 2,888 were classified as normal MFs. We found total agreement per object by all three experts in 2908 (78.17%) of the cases.
2.2 AtNorM-MD
The second dataset, Atypical and Normal Mitosis (AtNorM)-MultiDomain (MD) extends beyond human breast cancer and covers six domains, spanning both human and canine tumors. These include canine lung cancer, canine lymphoma, canine cutaneous mast cell tumor, human neuroendocrine tumor, canine soft tissue sarcoma, and human melanoma. It was sourced from all but the human breast cancer domains of the publicly available MIDOG++ dataset (Aubreville et al., 2023c). It comprises 2,107 mitotic figures, with 219 (10.4%) being atypical. Labeling was performed as a majority vote of five pathology experts, three of which were board-certified. We found total agreement per object by all five experts in 1,466 (69.58%) of the cases.
2.3 AtNorM-Br
The third dataset, AtNorM-Br, also made available publicly within the scope of this work, contains 746 MF instances from 179 patients from the breast cancer (BRCA) cohort of the The Cancer Genome Atlas (TCGA) (Lingle et al., 2016). TCGA contains images from various sources and with partially mixed quality and is thus also a valuable asset for the assessment of generalization. This dataset was annotated by a single expert with high experience in AMF classification, according to which 128 MFs were found to be atypical and the remaining 618 were classified as normal, yielding an AMF rate of 17.16%.
Table 2: Overview of foundation models used in this paper
Model Pretraining Algorithm Training Dataset Size Model Type Size (Params) WSIs Tiles UNI DINOv2 100K 100M ViT-L/16 307M UNI2-h DINOv2 350K 200M ViT-H/14 681M Virchow DINOv2 1.5M 2B ViT-H/14 632M Virchow2 DINOv2 3.4M 1.7B ViT-H/14 632M Prov-Gigapath DINOv2 170K 1.3B ViT-g/16 1.1B H-Optimus-0 DINOv2 500K - ViT-g/14 1.1B H-Optimus-1 DINOv2 1M 2B ViT-g/14 1.13B
3 Methods
3.1 End-to-end-trained Deep Learning Models
To establish robust baselines for atypical vs. normal mitoses classification, we evaluate three widely-used deep learning architectures – EfficientNetV2 (Tan and Le, 2021), ViT (Dosovitskiy et al., 2020), and Swin Transformer (Liu et al., 2021). All of these models have demonstrated strong performance across various medical imaging datasets and tasks. EfficientNetV2 is a well-established convolutional neural network (CNN) known for its strong performance and efficiency in various classification tasks. It serves as a representative of traditional CNN-based approaches. In contrast, the ViT represents a paradigm shift from convolutional architectures to transformer-based models. A ViT processes images as sequences of patches and leverages self-attention mechanisms to model global relationships across the input. This global receptive field allows the ViT to capture long-range dependencies, which can be particularly useful in complex histopathological patterns. The Swin Transformer further advances the transformer-based approach by introducing a hierarchical architecture with shifted windows. This design enables the model to compute attention locally while gradually building up to global representations, providing a balance between computational efficiency and the ability to capture both local details and global structure. Swin Transformers have been particularly effective in dense prediction tasks and high-resolution image analysis, making it well-suited for mitotic figure classification (Liu et al., 2021). Moreover, the selected model architectures were also chosen to be of similar size. Each model was trained in an end-to-end manner and evaluated on three held-out test datasets: a random, case-based split of the AMi-Br dataset, accounting for $\approx$ 22 % of the samples in the dataset (samples not used in training or validation), and the entire AtNorM-MD and AtNorM-Br datasets.
3.2 Foundation Models
Foundation models are models that are trained on large and diverse corpora of data, often using unsupervised techniques like self supervision. The goal of training is to train feature extractors that generalize well and are easily adaptable to down-stream tasks. In the field of computational pathology, we have seen a very strong incline in publicly available models, warranting an analysis on the task of AMF classification. We compare eight state-of-the-art models – UNI (Chen et al., 2024), UNI2-h (Chen et al., 2024; MahmoodLab, 2025), Virchow (Vorontsov et al., 2024), Virchow2 (Zimmermann et al., 2024), Prov-Gigapath (Xu et al., 2024), H-Optimus-0 (Saillard et al., 2024), H-Optimus-1 (Bioptimus, 2025), and H0-mini (Filiot et al., 2025). An overview of the details for each model is provided in Table 2. All of the selected foundation models are ViT-based and have been pre-trained on datasets comprising of hundreds of thousands to multiple millions of whole slide images.
3.2.1 Linear Probing of Foundation Models
We use a linear probing strategy, where the feature extractor (i.e., the foundation model) is used to extract the embeddings. After that, the foundation model is kept frozen, and a linear classifier is trained on top of the extracted features. This process helps us to assess the ability of the representations of the foundation models to classify atypical vs. normal mitoses.
3.2.2 LoRA Fine-Tuning
Parameter-Efficient Fine-Tuning (PEFT) methods (Houlsby et al., 2019) are model adaptation strategies which focus on fine-tuning models with minimal changes to their parameters. Traditional fine-tuning updates all model parameters, which becomes computationally expensive for large models. LoRA (Hu et al., 2022), a PEFT method, addresses this by freezing the original weights and introducing small, trainable low-rank matrices into specific layers, typically within the query, key, and value weight matrices of self-attention blocks. Only these matrices are updated during training, greatly reducing memory and compute requirements while retaining performance comparable to full fine-tuning. We apply LoRA-based finetuning to the same set of models previously evaluated with linear probing – UNI, UNI2-h, Virchow, Virchow2, Prov-Gigapath, H-Optimus-0 and H-Optimus-1, for a direct comparison.
[…]
4 Experimental Setup
We set up detailed experiments for the above cases. The image preprocessing for the binary classification experiments included resizing all patches to 224×224 pixels. Training augmentations consisted of random horizontal flip, rotation, color jitter, and random resized crop, followed by normalization with ImageNet statistics. Validation images were only resized and normalized. The best checkpoint for each fold was selected based on validation balanced accuracy. All experiments were performed using 5-fold cross validation, where stratification of the data was based on the slide (patient) to ensure that all patches from a given slide appeared in only one split (train or validation), thus preventing data leakage. To mitigate class imbalance, we employed weighted random sampling during training, with sample weights inversely proportional to class frequencies. All models were trained using the Adam optimizer with L2 regularization (learning rate: 1e-4, weight decay: 1e-5) and cross-entropy loss. We applied early stopping with a patience of 15 epochs based on validation balanced accuracy, and used learning rate scheduling on plateau (factor: 0.5, patience: 3, minimum LR: 1e-7). Each model was trained for up to 100 epochs with a batch size of 8. The rest of the case-specific details are described below in the subsections.
4.1 End-to-end-trained Deep Learning Models
We initialized all three models that were trained in an end-to-end fashion (i.e., EfficientNetV2, ViT and Swin Transformer) with ImageNet-pretrained weights, and the original classification head was replaced with a custom binary classifier consisting of a single linear layer. Specifically, we used the vit_large_patch16_224 model from the timm library, a ViT with 224×224 input resolution, 16×16 patch size, and 86.6 million parameters. The Swin Transformer variant used was swin_base_patch4_window7_224 which features a 4×4 patch size and 7×7 shifted attention windows, comprising 87.8 million parameters. For EfficientNetV2, we employed efficientnetv2_m from timm, which has approximately 54.1 million parameters. This protocol was applied identically to all three models to enable fair comparison of their performance on the atypical mitosis classification task.
4.2 Linear Probing with Foundation Models
We conducted comprehensive evaluation of seven foundation models using a standardized linear probing method on the AMi-Br dataset for atypical versus normal mitosis classification. For each model, we extracted high-dimensional feature embeddings from histopathology images and trained a single linear classification layer while keeping the pre-trained feature extractor frozen. This method was consistently applied across all six foundation models to ensure fair comparison of their feature representation capabilities on the atypical classification task.
4.3 LoRA Fine-tuning of Foundation Models
For the LoRA-based fine-tuning of the foundation models, each model was initialized with publicly available pre-trained weights and adapted using LoRA with rank 8, scaling factor 16, and dropout rate 0.3, applied to transformer attention and MLP layers. The classification head was re-initialized and trained jointly with LoRA modules, while the rest of the backbone remained frozen. This standardized protocol ensured fair and rigorous evaluation of LoRA-based adaptation across diverse foundation models on the AMi-Br benchmark.
Table 3: Performance of different models across three test datasets. Values are reported as mean $\pm$ standard deviation.
Model AMi-Br AtNorM-Br AtNorM-MD Balanced Acc. AUROC Balanced Acc. AUROC Balanced Acc. AUROC EfficientNetV2 $0.6488\pm 0.1265$ $0.7124\pm 0.1608$ $0.6457\pm 0.1055$ $0.7055\pm 0.1250$ $0.6034\pm 0.1166$ $0.6721\pm 0.1279$ Vision Transformer $0.7552\pm 0.0191$ $0.8634\pm 0.0104$ $0.7570\pm 0.0171$ $0.8478\pm 0.0189$ $0.7439\pm 0.0149$ $0.8396\pm 0.0213$ Swin Transformer $0.7491\pm 0.1061$ $0.8451\pm 0.0996$ $0.7213\pm 0.1116$ $0.8272\pm 0.0879$ $0.7226\pm 0.1048$ $0.8069\pm 0.1213$ UNI $0.6477\pm 0.0108$ $0.7171\pm 0.0080$ $0.6217\pm 0.0015$ $0.6774\pm 0.0136$ $0.5715\pm 0.0153$ $0.6241\pm 0.0222$ UNI2-h $0.6472\pm 0.0138$ $0.7087\pm 0.0107$ $0.6612\pm 0.0169$ $0.7296\pm 0.0115$ $0.5919\pm 0.0162$ $0.6297\pm 0.0173$ Virchow $0.6452\pm 0.0139$ $0.7010\pm 0.0077$ $0.6441\pm 0.0212$ $0.7037\pm 0.0107$ $0.5758\pm 0.0172$ $0.6175\pm 0.0227$ Virchow2 $0.6403\pm 0.0079$ $0.7040\pm 0.0052$ $0.6667\pm 0.0133$ $0.7201\pm 0.0060$ $0.5652\pm 0.0064$ $0.5878\pm 0.0211$ Prov-Gigapath $0.6091\pm 0.0140$ $0.6569\pm 0.0132$ $0.6005\pm 0.0091$ $0.6323\pm 0.0125$ $0.5601\pm 0.0153$ $0.5818\pm 0.0129$ H-Optimus-0 $0.6372\pm 0.0164$ $0.6926\pm 0.0062$ $0.6122\pm 0.0185$ $0.6903\pm 0.0217$ $0.5603\pm 0.0400$ $0.5935\pm 0.0423$ H-Optimus-1 $0.5918\pm 0.0165$ $0.6494\pm 0.0108$ $0.6139\pm 0.0195$ $0.6696\pm 0.0193$ $0.5915\pm 0.0308$ $0.6436\pm 0.0499$ UNI (LoRA) $0.7952\pm 0.0092$ $0.8839\pm 0.0059$ $0.7183\pm 0.0226$ $0.7979\pm 0.0142$ $0.7069\pm 0.0278$ $0.8222\pm 0.0224$ UNI2-h (LoRA) $0.7138\pm 0.0121$ $0.8153\pm 0.0211$ $0.7301\pm 0.0152$ $0.8228\pm 0.0143$ $0.6914\pm 0.0321$ $0.7616\pm 0.0415$ Virchow (LoRA) $0.7878\pm 0.0250$ $0.8891\pm 0.0150$ $\mathbf{0.7696}\pm 0.0198$ $0.8540\pm 0.0200$ $\mathbf{0.7705}\pm 0.0287$ $\mathbf{0.8641}\pm 0.0247$ Virchow2 (LoRA) $\mathbf{0.8135}\pm 0.0145$ $\mathbf{0.9026}\pm 0.0051$ $0.7632\pm 0.0190$ $\mathbf{0.8579}\pm 0.0117$ $0.7424\pm 0.0305$ $0.8503\pm 0.0171$ Prov-Gigapath (LoRA) $0.7602\pm 0.0113$ $0.8682\pm 0.0122$ $0.7263\pm 0.0296$ $0.8077\pm 0.0184$ $0.7007\pm 0.0228$ $0.8073\pm 0.0259$ H-Optimus-0 (LoRA) $0.5888\pm 0.0766$ $0.6159\pm 0.1276$ $0.5901\pm 0.0827$ $0.6224\pm 0.1220$ $0.5406\pm 0.0619$ $0.5826\pm 0.1263$ H-Optimus-1 (LoRA) $0.5908\pm 0.0909$ $0.6362\pm 0.1143$ $0.5699\pm 0.0752$ $0.6352\pm 0.1091$ $0.5617\pm 0.0874$ $0.6019\pm 0.1312$
5 Results
On the AMi-Br dataset, the best-performing model was Virchow2 with LoRA finetuning, achieving an average balanced accuracy of 0.8135 and an AUROC of 0.9026, outperforming all other models. Other LoRA-fine-tuned models such as UNI (0.7952 mean balanced accuracy) and Virchow (0.7878 balanced accuracy) also demonstrated strong performance, significantly improving over their respective linear probing counterparts. Among the end-to-end trained baseline models, the ViT achieved the highest balanced accuracy (0.7552) and AUROC (0.8634), outperforming both EfficientNetV2 and Swin Transformer. Foundation models with just linear probing lagged behind, with balanced accuracies around 0.59-0.64.
On the AtNorM-Br dataset, the trend was similar. Virchow with LoRA finetuning achieved the highest balanced accuracy of 0.7696 while Virchow2 with LoRA finetuning achieved the highest AUROC of 0.8579, followed by Virchow2 with LoRA finetuning (balanced accuracy: 0.7632). Other LoRA-fine-tuned models such as UNI2-h (0.7301) and UNI (0.7183) also showed consistent improvements over their linear probing variants. Among the baselines, the Vision Transformer maintained strong performance (balanced accuracy: 0.7570, AUROC: 0.8478), again outperforming EfficientNetV2 and Swin Transformer. Foundation models without LoRA adaptation performed moderately, with balanced accuracies between 0.60 and 0.66.
The AtNorM-MD dataset, which represents a distinct distribution shift, saw a general drop in performance across models. Virchow pretrained with LoRA performed the best, with a balanced accuracy of 0.7705 and AUROC of 0.8641. Other strong LoRA-fine-tuned models included Virchow2 (0.7424 balanced accuracy) and UNI (0.7069 balanced accuracy), all of which substantially outperformed their linear probing counterparts. The best baseline in this setting was again the Vision Transformer, achieving 0.7439 balanced accuracy and 0.8396 AUROC. In contrast, foundation models without LoRA adaptation performed less robustly. Across the board, LoRA fine-tuning consistently enhanced generalization performance under domain shift.
Across all the datasets, two of the models, H-Optimus-0 and H-Optimus-1, both fine-tuned with LoRA, consistently performed poorly, even less than their linear probing counterparts in balanced accuracy.
6 Discussion
Our evaluation confirmed that all foundation models exhibited some degree of generalization to the AMF identification task, but the extent of their effectiveness varied considerably. The application of LoRA fine-tuning revealed particularly striking performance differences across models.
The poor performance of H-Optimus-0 and H-Optimus-1 could be the result of a resolution mismatch, as the model expects input images trained at 0.5 microns/pixel resolution, whereas our dataset has images at 0.25 microns/pixel. In contrast, Virchow2 demonstrated superior performance, likely benefiting from training on higher resolution images using a 40x objective, making it more compatible with our dataset characteristics.
We also notice considerable drop in performance for the foundation models (both linear probing and LoRAs-fine-tuned) in the two external test datasets, as reflected by the balanced accuracy and AUROC scores. We attribute this to the domain shift between the datasets. For the AtNorM-Br dataset, the observed performance may partially also result from individually different criteria in the discrimination of AMFs, an effect that evens out when multiple raters are involved. As only a single annotator performed the labeling process for AtNorM-Br, this introduces potentially systematic label biases. Moreover, an image distribution shift cannot be ruled out, as the images were retrieved from different sources (labs, scanners, etc.), making the dataset more heterogeneous. For the AtNorM-MD dataset, where five annotators were involved, a systematic label bias is rather unlikely, however, the inclusion of multiple domains beyond breast cancer introduces a more pronounced image domain shift, encompassing a wider range of tissue types that pose significant challenges for model adaptation.
A significant limitation of our approach lies in training exclusively on the AMi-Br dataset with a relatively limited number of cases of only human breast cancer, restricting the models’ ability to learn the full spectrum of AMF representations across different tissue types, tumors, institutions, staining protocols, and scanners. Our results emphasize that automated AMF identification remains a highly challenging problem in computational pathology, requiring further methodological advances. Our new, multi-domain datasets, provided by this work, help further this cause.
EMA
exponential moving average
H&E
Hematoxylin & Eosin
RoI
region of interest
WSI
whole slide image
MF
mitotic figure
TCGA
The Cancer Genome Atlas
CMC
canine mammary carcinoma
AtNorM
Atypical and Normal Mitosis
AMF
atypical mitotic figure
CNN
convolutional neural network
ViT
vision transformer
LoRA
Low Rank Adaptation
ROC
receiver operating characteristic
AUROC
Area under the Receiver Operating Characteristic Curve
[…]
CAB, VW, and CK acknowledge the support from the Austrian Research Fund (FWF, project number: I 6555). SB, TC, RK, and MA acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number: 520330054). KB acknowledge support by the German Research Foundation (DFG) project 460333672 CRC1540 EBM.
[…]
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
[…]
We declare no conflicts of interest.
[…]
All data used in this study is public and can be found at this github repository: https://github.com/DeepMicroscopy/AMi-Br_Benchmark.